Simplify
(2x^-6y^-4z^0/3^-2x^-2y^-8z^5)^3

Answer by stanbon(57328)

Simplify
[(2 x^-6 y^-4 z^0)/(3^-2 x^-2 y^-8 z^5)]^3
=( 2^3 x^-18 y^-12 z^0) / (3^-6 x^-6 y^-24 z^15)
=(2^3 3^6)(x^6/x^18 )(y^24/y^12)(z^0/z^15)
=(5832) (1/x^12)(y^12)(1/z^15)
=5832[y^12/[x^12 z^15]
Cheers,
Stan H.